# Holographic metals at finite volume

### v1

@@ Abstract
We construct the electron star solution in asymptotically global AdS spacetime, and investigate its stability properties, both locally under perturbations and globally with respect to the Reissner-Nordström black hole and thermal AdS metrics. We interpret the resulting phase diagram as that of a holographic metal confined to a finite volume. We identify a quantum critical point at finite chemical potential, around which the different phases are organized.
@@ 2 The model
Notice that we are working not only in the limit $mL\gg 1$ but also in the classical limit $\kappa^{2}/L^{2}\ll 1$, and they are both in the definition of $\gamma$. However, as discuss in Hartnoll:2011 , we are interested in the scaled limit where $\gamma\sim 1$. This is obtained by taking $e^{-1}\sim\kappa/L\ll 1$, i.e. we require the gravitational coupling to be the square of the Maxwell coupling (‘probe brane’ limit). Furthermore, these limits implies for the scaled mass $\tilde{m}^{2}\sim\sqrt{\gamma}\sim 1$.
@@
where $T_{0}$ and $\mu_{0}$ are reference valus Note that with $h=0$, we recover the already studied holographic neutron star conditions Grandi:2018 ; Grandi:2020 .
We will solve now the equations of motions (10), satisfying the equations of state (16) in thermodynamic equilibrium (18). This equations of state are local and depend on $r$, thus we need to set some initial conditions at $r=0$. We will star the construction imposing null conditions at the origin (which will leads to a BH), and then with some central values (a pure electron star configuration).
@@ 3 Solutions: black hole and electron star
Beyond the horizon, one might think a possible configuration consisting of a cloud of charged particles in equilibrium surrounding the black hole. This type of star configuration has been studied in Hartnoll:2011 ; Thorlacius:2011 , where the temperature of the particle cloud is fixed at $T=0$. Consequently, these constructions are not in thermodynamic equilibrium with the black hole. If we consider the cloud at a non-zero temperature and impose thermodynamic equilibrium with the black hole, we find that these configurations are not stable (at least at the classical level), as explained in Appendix C. In conclusion, for a thermal star with a black hole at the origin, the only thermodynamically stable configuration arises when all particles eventually fall into the black hole, resulting in a pure Reissner-Nordström black hole solution.
@@
We thus construct the star profiles in terms of $T_{\infty}$ and $\mu_{\infty}$ by using the shooting method to map the central values $T_{0}$ and $\mu_{0}$ to the boundary, as shown in Fig. 2. As evident from the figure, these profiles correspond exclusively to ‘pure core’ stars: we observe a region characterized by an ‘abrupt edge’. Subsequently, a power-law behavior emerges and evolves into a ‘cusp’ star. The appearance of the power law denotes a critical region where the star solution becomes unstable. Consequently, a stable solution exists only within the region delimited by the critical curve. As we increase the temperature or chemical potential beyond this boundary, the star collapses. The critical curve is constructed using the Katz criterion, where each critical point is identified by the vertical asymptote where the first derivative of the curve changes sign Katz:1978 ; Katz:2002  [^2] The Katz criterion identifies a vertical asymptote in the caloric curve $M_{s}(T_{\infty}^{-1})$ while keeping $\mu_{\infty}/T_{\infty}$ fixed. This constraint ensures that we compare star solutions with the same entropy per particle $s=S/N$. In the neutron star case Canavesi:2023 , this ratio is constant throughout the bulk, so fixing the central parameters $\mu_{0}/T_{0}$ is equivalent to fixing the boundary values. In the present electron star case, however, the ratio varies radially. From (23), we have $\mu_{\infty}/T_{\infty}=(\mu_{0}+h_{\infty})/T_{0}$, where $h_{\infty}$ is obtained numerically. Nevertheless, since $h_{\infty}$ remains small up to the critical region (see Fig. 2), we approximate the Katz curves by fixing $\mu_{0}/T_{0}$. This approximation introduces a discrepancy of $\sim 15\%$ between the critical central and boundary values. This instability signals a second-order phase transition, which would occur well after the first-order transition discussed in Sec. 4..
@@
Figure 2: Top: (Left) The shaded region indicates the domain where the Electron Star is a stable solution according to the Katz criterion. This region is delimited by the red critical curve, which is constructed by varying the initial values $\mu_{0}$ and $T_{0}$ and identifying the critical point on the corresponding Katz curve. The colored curves represent families of solutions with fixed $\mu_{0}/T_{0}$, where circles and squares denote representative stable and unstable configurations, respectively. The solid black line marks the star-black hole phase transition derived in Sec. 4. (Right) Katz stability curves for the families indicated in the phase diagram. The vertical dashed red lines mark the critical point obtained for each curve. Bottom: Log-Log plor for the radial density profiles corresponding to the specific configurations marked in the top-left panel; the colors match the phase space trajectories. Solid lines represent stable solutions (circles in the top-left figure), characterized by an abrupt edge profile, while dashed lines correspond to unstable solutions (squares) exhibiting power-law behavior.
@@
Having identified the domain where the electron star solution exists, we must now determine its thermodynamic stability relative to the vacuum solutions, namely the Reissner-Nordström black hole and TAdS. To this end, in the following section, we will evaluate and compare the free energies of these geometries to construct the complete phase diagram of the system.
@@ 4 Phase Transitions
The details of each Euclidean on-shell action computation are provided in Appendix B. Notice that since the action is of order $1/\kappa^{2}$ (which is large in the classical limit), the partition function is exponentially dominated by the term with the smallest action (i.e. the minimum free energy). As explained in the Appendix, we defined the temperature and chemical potential in the previous section ensuring that all three saddle points share the same boundary values $T_{\infty}$ and $\mu_{\infty}$. Thus, by expressing the actions in terms of $T_{\infty}$ and $\mu_{\infty}$, we map the phase space and determine the dominant phase. The resulting free energies to be compared are given by
@@
Figure 3: Left: Phase diagram of the holographic electron star in the boundary plane $(\mu_{\infty},T_{\infty})$. The colored region indicates where the electron star is the dominant solution (minimal free energy), while the blue line and grey region correspond to TAdS and black hole dominance, respectively. The solid black lines denote first-order phase transition, and the dashed black line indicates where the Hawking-Page transition would occur in the absence of the star Chamblin:1999 . The colored lines (green, brown, red) represent the slices along which we evaluate the free energies shown in the right panels. Right: Free energy comparisons for the electron star (orange), black hole (grey), and TAdS (blue) corresponding to the cuts in the phase diagram. The frame colors match the cuts in the left panel. The black dots mark the phase transition points. The red frame corresponds to the $T_{\infty}=0$ case, from which the quantum critical point is determined.
@@ 5 Discussion and future directions
We have constructed the spherically symmetric holographic electron star solution at finite temperature $T$ and chemical potential $\mu$. We performed a Katz equilibrium analysis and found that the system possesses a stable branch that eventually destabilizes at a certain temperature, never regaining stability. This implies the existence of a finite region in the boundary phase space $(\mu_{\infty},T_{\infty})$ where the star is stable, extending up to a critical boundary where a second-order phase transition occurs. Contrary to the neutron star case previously studied in Grandi:2018 , the electron star only exhibits a dense core structure when it is stable, while the power-law behavior emerges strictly when the solution becomes unstable.
@@ Appendix A Details of the model
We aim to describe the thermodynamics of a large number of charged self-gravitating fermions in equilibrium (also known as ‘Electron Stars’) within a holographic framework. To this end, we consider a 3+1 dimensional global AdS spacetime and approximate the matter dynamics as a charged perfect fluid coupled to the gravitational field. The action functional consists of three contributions: the Einstein-Hilbert ($S_{\text{Eins.}}$), Maxwell ($S_{\text{Mxwl.}}$), and ideal fluid ($S_{\text{Fluid}}$) terms
$S=\int d^{4}x\>\sqrt{-g}(\mathcal{L}_{Eins.}+\mathcal{L}_{Mxwl.}+\mathcal{L}_{Fluid})$ (31)
@@
$\mathcal{L}_{Eins.}$ $=$ $\frac{1}{2\kappa^{2}}\left(R+\frac{6}{L^{2}}\right)$ (32)
$\mathcal{L}_{Mxwl.}$ $=$ $-\frac{1}{4e^{2}}F_{\mu\nu}F^{\mu\nu}$ (33)
$\mathcal{L}_{Fluid}$ $=$ $-\rho(\sigma,s)+\sigma u^{\mu}(\partial_{\mu}\phi+\theta\>\partial_{\mu}s+A_{\mu})+\lambda(u^{\mu}u_{\mu}+1)$ (34)
@@ Appendix B Free energy computation details
$I=-\int d^{4}x\>\sqrt{g}\,\mathcal{L}_{\sf bulk}+I_{\sf GHY}+I_{\sf ct}$ (53)
where $I_{\sf ct}$ is the counterterm action required for asymptotically $AdS_{4}$ spaces Balasubramanian:1999 ,
@@ Appendix C Black hole surrounded by an electron cloud
where $H(\epsilon)$ represents the density of states function appearing in the definitions of $\rho$, $\sigma$, and $P$ in (16).
If we attempt to construct a thermodynamically stable solution for a cloud extending to the black hole horizon, a divergence arises. Specifically, performing a Sommerfeld expansion at low temperatures (see Appendix E in Pathria:2011 ), we obtain
$\mathcal{I}\approx\int_{1}^{\mu}\,H(\epsilon)\>d\epsilon+\frac{\pi^{2}}{6}\,T^{2}\,H^{\prime}(\mu)+\mathcal{O}(T^{4})$ (83)
The first term corresponds to the zero-temperature limit (fully degenerate cloud). The issue arises with the finite-temperature corrections: if we impose equilibrium with the black hole, $T_{0}=T_{\text{\tiny BH}}$, then the local temperature $T(r)$ diverges as $r\to r_{0}$ (since $f(r_{0})=0$). Consequently, the fluid density and pressure would diverge at the horizon. The only consistent solution in this setup is to set $T_{0}=0$, i.e., the electron cloud must be at zero temperature. In this scenario, the fluid density vanishes at the horizon and the cloud only exists in regions where the local chemical potential satisfies $\mu(r)\geq 1$. This results in a black hole surrounded by a zero-temperature cloud, potentially with a global temperature parameter shifted by gravitational backreaction, but not in thermal contact with the horizon Thorlacius:2011 .

### v2

@@ Abstract
We construct the electron star solution in asymptotically global AdS spacetime, and investigate its stability properties, both locally under perturbations and globally with respect to the Reissner-Nordström black hole and thermal AdS spacetimes. We interpret the resulting phase diagram as that of a holographic metal confined to a finite volume. We identify a quantum critical point at finite chemical potential, around which the different phases are organized.
@@ 2 The model
Notice that we are working not only in the limit $mL\gg 1$ but also in the classical limit $\kappa^{2}/L^{2}\ll 1$, and they are both in the definition of $\gamma$. However, as discussed in Hartnoll:2011 , we are interested in the scaled limit where $\gamma\sim 1$. This is obtained by taking $e^{-1}\sim\kappa/L\ll 1$, i.e. we require the gravitational coupling to be the square of the Maxwell coupling (‘probe brane’ limit). Furthermore, these limits implies for the scaled mass $\tilde{m}^{2}\sim\sqrt{\gamma}\sim 1$.
@@
where $T_{0}$ and $\mu_{0}$ are reference valus Note that when the electromagnetic field vanishes $h=0$, we recover the already studied holographic neutron star conditions Grandi:2018 ; Grandi:2020 .
We will solve now the equations of motions (10), satisfying the equations of state (16) in thermodynamic equilibrium (18). The equations of state are local and depend on $r$, thus we need to set some initial conditions at $r=0$. We start the construction imposing null conditions at the origin (which will leads to a BH), and then with some central values (a pure electron star configuration).
@@ 3 Solutions: black hole and electron star
Outside the horizon, one might think of a possible configuration consisting of a cloud of charged particles surrounding the black hole. This type of star configurations have been studied in Hartnoll:2011 ; Thorlacius:2011 . There, the temperature of the particle cloud was fixed to $T=0$, implying that the cloud is not in thermodynamic equilibrium with the black hole. If we consider the cloud at a non-zero temperature in order to impose thermodynamic equilibrium with the black hole, we find that these configurations are not stable (at least at the classical level), as explained in Appendix C. In conclusion, for a thermal star with a black hole around the origin, the only thermodynamically stable configuration arises when all particles eventually fall into the black hole, resulting in a pure Reissner-Nordström solution.
@@
We construct the star profiles in terms of $T_{\infty}$ and $\mu_{\infty}$ by shooting from the central values $T_{0}$ and $\mu_{0}$ to the boundary, as shown in Fig. 2. The critical curve is constructed using the Katz criterion Katz:1978 ; Katz:2002 . This criterion identifies a vertical asymptote in the caloric curve $M_{s}$ vs. $1/T_{\infty}$ while keeping $\mu_{\infty}/T_{\infty}$ fixed (this constraint ensures that we compare star configurations with the same entropy per particle $s=S/N$). In the holographic neutron star case Canavesi:2023 , this ratio is constant throughout the bulk, meaning that fixing the central parameters $\mu_{0}/T_{0}$ is equivalent to fixing the boundary values. In the present electron star case, however, the ratio varies radially, such that $\mu_{0}/T_{0}\neq\mu_{\infty}/T_{\infty}$ (as seen from (24)). To address this, we first generate a manifold of solutions by varying parameters while keeping the central ratio $\mu_{0}/T_{0}$ fixed. Using the map (24), we then translate these solutions into the boundary parameter space $M_{s}$ vs. $(1/T_{\infty},\mu_{\infty}/T_{\infty})$. Finally, to construct the proper caloric curves required by the Katz criterion, we extract contours of constant $\mu_{\infty}/T_{\infty}$ from this mapped surface. This procedure yields the Katz curves shown in Fig. 2. The critical boundary is then formed by the set of critical points, each identified by a vertical asymptote in the corresponding curve; the slight oscillations observed on the critical curve are numerical artifacts arising from the surface interpolation process.
@@
Figure 2: Top: (Left) The shaded region indicates the domain where the Electron Star solution exists. This region is delimited by the orange critical curve constructed by varying the initial values $\mu_{0}$ and $T_{0}$ and identifying the critical point on the corresponding Katz curve in the $\mu_{\infty}$ and $T_{\infty}$ space. The colored curves represent families of solutions with fixed $\mu_{\infty}/T_{\infty}$, where circles and squares denote representative stable and unstable configurations, respectively. The solid black line marks the star-black hole phase transition derived in Sec. 4. (Right) Katz stability curves for the families indicated in the phase diagram. The vertical dashed orange lines mark the critical point obtained for each curve, while the circles and squares denote the stable and unstable configurations shown below. Bottom: Log-Log plot for the radial density profiles corresponding to the specific configurations marked in the top-left panel; the colors match the phase space trajectories. Solid lines represent stable solutions (circles in the top figure), characterized by an abrupt edge profile, while dashed lines correspond to unstable solutions (squares) exhibiting power-law behavior.
As seen in Fig. 2, according to the Katz criterion, stable and unstable solutions coexist in the $\mu_{\infty}$ vs. $T_{\infty}$ plane only within the domain delimited by the “critical” (orange) curve, although they correspond to different central values; beyond that region, no electron star solutions exist. In terms of the central parameters, stable solutions correspond to low values of $\mu_{0}$ and $T_{0}$, while the unstable configurations emerge as these parameters are increased. Furthermore, the stable density profiles shown in the figure correspond exclusively to ‘pure core’ stars, which are characterized by an abrupt edge. Conversely, the unstable solutions exhibit a power-law behavior at the edge of the profile, eventually evolving into a ‘cusp’ star. This instability signals a second-order phase transition, which occurs well after the first-order transition that we will discuss in Sec. 4.
@@
Having established the electron star domain, we must now determine its thermodynamic stability relative to the vacuum solutions, namely the Reissner-Nordström black hole and TAdS. To this end, in the following section, we will evaluate and compare the free energies of these geometries to construct the complete phase diagram of the system.
@@ 4 Phase transitions
The details of each Euclidean on-shell action computation are provided in Appendix B. Notice that since the action is of order $1/\kappa^{2}$ (which is large in the classical limit), the partition function is exponentially dominated by the term with the smallest action (i.e. the minimum free energy). We have defined the temperature and chemical potential in the previous section ensuring that all three saddle points share the same boundary values $T_{\infty}$ and $\mu_{\infty}$. Thus, by expressing the on-shell actions in terms of $T_{\infty}$ and $\mu_{\infty}$, we map the phase space and determine the dominant phase. The resulting free energies to be compared are given by
@@
Figure 3: Left: Phase diagram of the holographic electron star in the boundary plane $(\mu_{\infty},T_{\infty})$. The colored region indicates where the electron star is the dominant solution (minimal free energy), while the red line and grey region correspond to TAdS and black hole dominance, respectively. The solid black lines denote first-order phase transition, and the dashed black line indicates where the Hawking-Page transition would occur in the absence of the star Chamblin:1999 . The colored lines (green, brown, blue) represent the slices along which we evaluate the free energies shown in the right panels. Right: Free energy comparisons for the electron star (blue), black hole (grey), and TAdS (red) corresponding to the cuts in the phase diagram. The frame colors match the cuts in the left panel. The black dots mark the phase transition points. The blue frame corresponds to the $T_{\infty}=0$ case, from which the quantum critical point is determined.
@@ 5 Discussion and future directions
We have constructed the spherically symmetric holographic electron star solution at finite temperature $T$ and chemical potential $\mu$. We performed a Katz equilibrium analysis and found that the system possesses a stable branch that eventually destabilizes at a certain central temperature, never regaining stability. This implies the existence of a finite region in the boundary phase space $(\mu_{\infty},T_{\infty})$ where the star is stable, extending up to a critical boundary where a second-order phase transition occurs. Contrary to the neutron star case previously studied in Grandi:2018 , the electron star only exhibits a dense core structure when it is stable, while the power-law behavior emerges strictly when the solution becomes unstable.
@@ Appendix A Details of the model
We aim to describe the thermodynamics of a large number of charged self-gravitating fermions in equilibrium (also known as ‘Electron Stars’) within a holographic framework. To this end, we consider a 3+1 dimensional global AdS spacetime and approximate the matter dynamics as a charged perfect fluid coupled to the gravitational field. The action functional consists of three contributions: the Einstein-Hilbert ($S_{\sf{Eins.}}$), Maxwell ($S_{\sf{Mxwl.}}$), and ideal fluid ($S_{\sf{Fluid}}$) terms
$S=\int d^{4}x\>\sqrt{-g}(\mathcal{L}_{\sf Eins.}+\mathcal{L}_{\sf Mxwl.}+\mathcal{L}_{\sf Fluid})$ (31)
@@
$\mathcal{L}_{\sf Eins.}$ $=$ $\frac{1}{2\kappa^{2}}\left(R+\frac{6}{L^{2}}\right)$ (32)
$\mathcal{L}_{\sf Mxwl.}$ $=$ $-\frac{1}{4e^{2}}F_{\mu\nu}F^{\mu\nu}$ (33)
$\mathcal{L}_{\sf Fluid}$ $=$ $-\rho(\sigma,s)+\sigma u^{\mu}(\partial_{\mu}\phi+\theta\>\partial_{\mu}s+A_{\mu})+\lambda(u^{\mu}u_{\mu}+1)$ (34)
@@ Appendix B Free energy computation details
$I=-\int d^{4}x\>\sqrt{g}\,(\mathcal{L}_{\sf Eins.}+\mathcal{L}_{\sf Mxwl.}+\mathcal{L}_{\sf Fluid})+I_{\sf GHY}+I_{\sf ct}$ (53)
where $I_{\sf ct}$ is the holographic renormalization counterterm in asymptotically $AdS_{4}$ spaces Balasubramanian:1999 ,
@@ Appendix C Black hole surrounded by an electron cloud
where $H(\epsilon)$ represents the “density of states” function appearing in the definitions of $\rho$, $\sigma$, and $P$ in (16).
Notice that attempting to construct a thermodynamically stable solution for a cloud extending down to the black hole horizon leads to a divergence. If we impose thermal equilibrium with the black hole by setting $T_{0}=T_{\text{\tiny BH}}$, the local temperature $T(r)$ diverges as $r\to r_{0}$ (since $f(r_{0})=0$). Consequently, the local fluid density and pressure would diverge at the horizon. The only consistent solution in this setup is to set $T_{0}=0$, i.e., the electron cloud must be at zero temperature (a fully degenerate Fermi gas). In this scenario, the fluid density vanishes at the horizon and the cloud only exists in regions where the local chemical potential satisfies $\mu(r)\geq 1$. This results in a black hole surrounded by a zero-temperature cloud, potentially with a global temperature parameter shifted by gravitational backreaction, but not in thermal contact with the horizon Thorlacius:2011 .